### v3

# Extreme Compression of Large Language Models via Additive Quantization

Abstract
The emergence of accurate open large language models (LLMs) has led to a race towards performant quantization techniques which can enable their execution on end-user devices. In this paper, we revisit the problem of “extreme” LLM compression—defined as targeting extremely low bit counts, such as 2 to 3 bits per parameter—from the point of view of classic methods in Multi-Codebook Quantization (MCQ). Our algorithm, called AQLM, generalizes the classic Additive Quantization (AQ) approach for information retrieval to advance the state-of-the-art in LLM compression, via two innovations: 1) learned additive quantization of weight matrices in input-adaptive fashion, and 2) joint optimization of codebook parameters across each transformer blocks. Broadly, AQLM is the first scheme that is Pareto optimal in terms of accuracy-vs-model-size when compressing to less than 3 bits per parameter, and significantly improves upon all known schemes in the extreme compression (2bit) regime. In addition, AQLM is practical: we provide fast GPU and CPU implementations of AQLM for token generation, which enable us to match or outperform optimized FP16 implementations for speed, while executing in a much smaller memory footprint.
Machine Learning, Quantization
1 Introduction
Figure 1: Comparison of AQLM (2-bit) relative to the state-of-the-art QuIP# (2-bit) and the original 16-bit weights on Llama 2 7, 13, and 70B models.
The rapid advancement of generative large language models (LLMs) has led to massive industrial and popular interest, driven in part by the availability of accurate open LLMs, such as Llama 1 and 2 (Touvron et al., 2023), Falcon (TII UAE, 2023), BLOOM (Scao et al., 2022), OPT (Zhang et al., 2022), or NeoX/Pythia (Biderman et al., 2023). A key advantage of open models is that they can be inferenced or fine-tuned locally by end-users, assuming that their computational and memory costs can be reduced to be manageable on commodity hardware. This has led to several methods for inference and fine-tuning on compressed LLMs (Dettmers et al., 2022; Frantar et al., 2022a; Dettmers & Zettlemoyer, 2022; Lin et al., 2023; Dettmers et al., 2023a). Currently, the primary approach for accurate post-training compression of LLMs is quantization, which reduces the bit-width at which model weights (and possibly activations) are stored, leading to improvements in model footprint and memory transfer.
By and large, LLM weights are compressed via “direct” quantization, in the sense that a suitable quantization grid and normalization are first chosen for each matrix sub-component, and then weights are each mapped onto the grid either by direct rounding, e.g. (Dettmers & Zettlemoyer, 2022), or via more complex allocations, e.g. (Frantar et al., 2022a). Quantization induces a natural compression-vs-accuracy trade-off, usually measured in terms of model size vs model perplexity (PPL). Existing approaches can achieve arguably low accuracy loss at 3-4 bits per element (Dettmers et al., 2023b; Chee et al., 2023; Kim et al., 2023), and can even stably compress models to 2 or even less bits per element, in particular, for extremely large models (Frantar & Alistarh, 2023). Yet, in most cases, low bit counts come at the cost of significant drops in accuracy, higher implementation complexity and runtime overheads. Specifically, from the practical perspective, “extreme” quantization in the 2-bit range using current techniques is inferior to simply using a smaller base model and quantizing it to higher bitwidths, such as 3-4 bits per parameter, as the latter yields higher accuracy given the same model size in bytes (Dettmers & Zettlemoyer, 2022; Chee et al., 2023).
Contribution. In this work, we improve the state-of-the-art in LLM compression by showing for the first time that Multi-Codebook Quantization (MCQ) techniques can be extended to LLM weight compression. Broadly, MCQ is a family of information retrieval methods (Chen et al., 2010; Jegou et al., 2010; Ge et al., 2013; Zhang et al., 2014; Babenko & Lempitsky, 2014; Martinez et al., 2016, 2018), consisting of specialized quantization algorithms to compress databases of vectors, allowing for efficient search. Unlike direct quantization, MCQ compresses multiple values jointly, by leveraging the mutual information of quantized values.
More precisely, we extend Additive Quantization (AQ) (Babenko & Lempitsky, 2014; Martinez et al., 2016), a popular MCQ algorithm, to the task of compressing LLM weights such that the output of each layer and Transformer block are approximately preserved. Our extension reformulates the classic AQ optimization problem to reduce the error in LLM layer outputs under the input token distribution and as well as to jointly optimize codes over layer blocks, rather than only preserving the weights themselves as in standard AQ. We refer to the resulting procedure as Additive Quantization of Language Models (AQLM). Unlike some extreme LLM quantization approaches that require hybrid sparse-quantized formats which separate outlier quantization (Kim et al., 2023; Dettmers et al., 2023b), AQLM quantizes models in a simple homogeneous format, which is easy to support in practice. Our main contributions are as follows:
  1. 1.
    We propose the AQLM algorithm, which extends AQ to post-training compression of LLM weights, via two innovations: (1) adapting the MAP-MRF [^1] Maximum a Posteriori inference in Markov Random Fields optimization problem behind AQ to be instance-aware, taking layer calibration input & output activations into account; (2) complementing the layer-wise optimization with an efficient intra-block tuning technique, which optimizes quantization parameters jointly over several layers, using only the calibration data.
  2. 2.
    We evaluate the effectiveness of this algorithm on the task of compressing accurate open LLMs from the Llama 2 (Touvron et al., 2023) family with compression rates of 2-4 bits per parameter. We find that AQLM outperforms the previous state-of-the-art across the standard 2-4 bit compression range, with the most significant improvements for extreme 2-bit quantization (see Figure 1). We provide detailed ablations for the impact of various algorithm parameters, such as code width and number of codebooks, and extend our analysis to the recent Mixtral model (Jiang et al., 2024). We also evaluate AQLM with improved fine-tuning algorithms from subsequent works, which leads to further increase in accuracy for 2- and 3-bit models.
  3. 3.
    We show that AQLM is practical, by providing efficient GPU and CPU kernels implementations for specific encodings, as well as end-to-end generation [^2] https://github.com/Vahe1994/AQLM. Results show that our approach can match or even outperform the floating point baseline in terms of speed, while reducing the memory footprint by up to 8x. Specifically, AQLM can be executed with layer-wise speedups of $\sim 30\%$ for GPUs, and of up to 4x for CPU inference.
2 Background & Related Work
2.1 LLM Quantization
Early efforts towards post-training quantization (PTQ) methods (Nagel et al., 2020; Gholami et al., 2021) that scale to LLMs such as ZeroQuant (Yao et al., 2022), LLM.int8() (Dettmers et al., 2022), and nuQmm (Park et al., 2022) employed direct round-to-nearest (RTN) projections, and adjusted quantization granularity to balance memory efficiency and accuracy. GPTQ (Frantar et al., 2022a) proposed a more accurate data-aware approach via an approximate large-scale solver for minimizing layer-wise $\ell_{2}$ errors.
Dettmers & Zettlemoyer (2022) examined the accuracy-compression trade-offs of these early methods, suggesting that 4-bit quantization may be optimal for RTN quantization, and observing that data-aware methods like GPTQ allow for higher compression, i.e. strictly below 4 bits/weight, maintaining Pareto optimality. Our work brings this Pareto frontier below 3 bits/weight, for the first time. Parallel work quantizing both weights and activations to 8-bits, by Dettmers et al. (2022)Xiao et al. (2022), and Yao et al. (2022) noted that the “outlier features” in large LLMs cause substantial errors, prompting various mitigation strategies.
Recently, several improved techniques have focused on the difficulty of quantizing weight outliers, which have high impact on the output error. SpQR (Dettmers et al., 2023b) addresses this by saving outliers as a highly-sparse higher-precision matrix. AWQ (Lin et al., 2023) reduces the error of quantizing channels with the highest activation magnitudes by employing per-channel scaling to reduce the error on important weights. SqueezeLLM (Kim et al., 2023) uses the diagonal Fisher as a proxy for the Hessian and implements non-uniform quantization through K-means clustering.
The published state-of-the-art method is QuIP (Chee et al., 2023). Concurrent to our work, an improved variant called QuIP# (Tseng et al., 2024) was introduced. Roughly, they work by first “smoothening” weights by multiplying with a rotation matrix, and then mapping them onto a lattice. At a high level, QuIP and QuIP# aim to minimize the “worst-case” error for each layer, given initial weights and calibration data. For instance, in QuIP#, the distribution of the rotated weights approximates a Gaussian, while the encoding lattice (E8P) is chosen to minimize “rounding” error. By contrast, our approach uses a different weight encoding (codebooks are additive), and learned codebooks instead of a fixed codebook. Thus, our insight is that we should be able to obtain higher accuracy by direct optimization of the codebooks over the calibration set, removing the rotation. Further, we show that codebooks for different layers can co-train via joint fine-tuning over the calibration data.
2.2 Quantization for Nearest Neighbor Search
Our work builds on approximate nearest neighbor search (ANN) algorithms. Unlike PTQ, ANN quantization aims to compress a database of vectors to allow a user to efficiently compute similarities and find nearest neighbors relative to a set of query points. For high compression, modern ANN search algorithms employ vector quantization (VQ)—which quantizes multiple vector dimensions jointly (Burton et al., 1983; Gray, 1984). It achieves this by learning “codebooks”: i.e. a set of learnable candidate vectors that can be used to encode the data. To encode a given database vector, VQ splits it into sub-groups of entries, then encodes every group by choosing a vector from the learned codebook. The algorithm efficiently computes distances or dot-products for similarity search by leveraging the linearity of dot products.
Quantization methods for ANN search generalize vector quantization and are referred to as multi-codebook quantization (MCQ). MCQ methods typically do not involve information loss on the query side, which makes them the leading approach for memory-efficient ANN (Ozan et al., 2016; Martinez et al., 2018). We briefly review MCQ below.
Product quantization (PQ) (Jegou et al., 2010) is an early version of MCQ, which encodes each vector $x\in\mathbf{R}^{D}$ as a concatenation of $M$ codewords from $M$ $\frac{D}{M}$-dimensional codebooks $C_{1},\ldots,C_{M}$, each containing $K$ codewords. PQ decomposes a vector into $M$ separate subvectors and applies vector quantization (VQ) to each subvector, while using a separate codebook. Thus, each vector $x$ is encoded by a tuple of codeword indices $[i_{1},\ldots,i_{M}]$ and approximated by $x\approx[c_{1i_{1}},\ldots,c_{Mi_{M}}]$. Fast Euclidean distance computation becomes possible using lookup tables:
$\|q-x\|^{2}\approx\|q-[c_{1i_{1}},\ldots,c_{Mi_{M}}]\|^{2}=\sum\limits_{m=1}^{M}{\|q_{m}-c_{mi_{m}}\|}^{2},$
where $q_{m}$ is the $m$th subvector of a query $q$. This sum can be calculated using $M$ additions and lookups if the distances from query subvectors to codewords are precomputed. Since product-based approximations work better if the $\frac{D}{M}$-dimensional components independent distributions, subsequent work has looked into finding better transformations (Ge et al., 2013; Norouzi & Fleet, 2013). As for the other similarity functions, (Guo et al., 2016) proposes a quantization procedure for maximum inner product search (MIPS). They minimize quantization error in the inner products between database and query vectors by solving a constrained optimization problem. Similarly to the formula above, this procedure allows for efficient inner product search by precomputing dot products between the query $q$ an all codes in the learned codebooks, then adding these partial dot products to recover the full similarity score.
Non-orthogonal quantizations. Follow-up work (Chen et al., 2010; Babenko & Lempitsky, 2014; Martinez et al., 2016; Zhang et al., 2014; Ozan et al., 2016; Martinez et al., 2018) generalized the idea of Product Quantization by approximating each vector by a sum of $M$ codewords instead of concatenation. The resulting procedure is still efficient while the approximation accuracy is increased.
For this, Residual Vector Quantization (Chen et al., 2010), quantizes original vectors, and then iteratively quantizes the approximation residuals from the previous iteration. Additive Quantization (AQ) (Babenko & Lempitsky, 2014) is more general, as it does not impose constraints on the codewords from the different codebooks. Usually, AQ provides the smallest compression errors, but is more complex to train for large $M$. We discuss this in detail in Section 3.
Finally, several recent works (Martinez et al., 2016, 2018; Zhang et al., 2014) elaborate the idea of Additive Quantization, proposing the more effective procedure for codebooks learning. Composite Quantization (CQ) (Zhang et al., 2014) learns codebooks with a fixed value of inner product between the codewords from different codebooks. Currently, the state-of-the-art compression accuracy is achieved by the LSQ method (Martinez et al., 2018).
Vector quantization for model compression.
There has been significant work on exploiting vector quantization in the context of machine learning. For instance, Zhou et al. (2017); Li et al. (2017); Chen et al. (2019) use multi-codebook quantization to compress word embeddings within deep learning models. Another line of work (Blalock & Guttag, 2021; McCarter & Dronen, 2022; Fernández-Marqués et al., 2023) explores vector quantization for linear models, or linear layers within deep models. Similarly to PQ above, these techniques pre-compute inner products between inputs and all codes, then compute linear layer via look-up, which speeds up inference. However, these algorithms introduce significant prediction error that does not allow them to compress deep models. Thus, we believe we are the first to successfully adapt and scale MCQ to LLMs.
3 AQLM: Additive Quantization for LLMs
3.1 Overview
We start from the observation that additive quantization (AQ) solves a related problem to post-training quantization (PTQ) (Nagel et al., 2020; Frantar et al., 2022b): both settings assume the existence of a set of “input” vectors, i.e. input data for AQ, and the weight matrix rows for PTQ. The goal is to compress these inputs while preserving dot product similarity, against query vectors (for AQ), and against layer input embeddings (for PTQ). The difference between the two is that AQ assumes that the distribution of queries is unknown, whereas PTQ methods, e.g. (Frantar et al., 2022b), show that it is sufficient to optimize for sample input embeddings from a set of calibration data.
At a high level, we start by solving the following problem: for a linear layer with $d_{in}$ input and $d_{out}$ output features given its weights $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ and a set of calibration inputs $\mathbf{X}\in\mathbb{R}^{d_{in}\times n}$, one seeks for a configuration of quantized weights $\hat{W}$ that optimizes squared error between the output of the original and compressed layer:
$\underset{\widehat{\mathbf{W}}}{\arg\min}||\mathbf{W}\mathbf{X}-\widehat{\mathbf{W}}\mathbf{X}||_{2}^{2}.$ (1)
In the following, we will assume that $\widehat{\mathbf{W}}$ is quantized using AQ, and adopt standard notation (Martinez et al., 2016). AQ splits weight rows into groups of $g$ consecutive elements, and represents each group of weights as a sum of $M$ vectors chosen from multiple learned codebooks $C_{1},...,C_{M}$, each containing $2^{B}$ vectors (for $B$-bit codes). A weight is encoded by choosing a single code from each codebook and summing them up. We denote this choice as a one-hot vector $b_{m}$, which results in the following representation for a group: $\sum_{m=1}^{M}C_{m}b_{ijm}$. This is similar to PTQ algorithms (Frantar et al., 2022a), except for using much more complex coding per group. To represent the full weights, we simply concatenate:
$\widehat{\mathbf{W}}_{i}{=}\sum_{m=1}^{M}C_{m}b_{i,1,m}\oplus...\oplus\sum_{m=1}^{M}C_{m}b_{i,d_{in}/g,m},$ (2)
where $\oplus$ denotes concatenation and $b_{ijm}\in\mathbb{R}^{2^{B}}$ represents a one-hot code for the $i$-th output unit, $j$-th group of input dimensions and $m$-th codebook.
Figure 2: Groups of weights are represented by a sum of codes selected from codebooks by corresponding indices.
Our algorithm will learn codebooks $C_{m}\in\mathbb{R}^{g\times 2^{B}}$ and the discrete codes represented by one-hot $b\in\mathbb{R}^{d_{out}\times d_{in}/g\times M\times 2^{B}}$. The resulting scheme encodes each group of $g$ weights using $M\cdot B$ bits and further requires $g\cdot 2^{B}\cdot 16$ bits for FP16 codebooks. The error becomes:
$\underset{C,b}{\arg\min}||\mathbf{W}\mathbf{X}-\left(\text{Concat}_{i,j}\sum^{M}_{m=1}C_{m}b_{i,j,m}\right)\mathbf{X}||_{2}^{2}.$ (3)
To learn this weight representation, we initialize codebooks $C$ and codes $b$ by running residual K-means as in Chen et al. (2010). Specifically, the initialization algorithm proceeds as follows: first, it runs K-means clustering of weight groups and saves the resulting cluster indices. Next, it computes the quantization errors by subtracting the nearest cluster from every weight. Finally, the algorithm runs another round of K-means clustering, but this time on quantization errors instead of weights. Thus, each subsequent codebook is initialized to compensate the quantization error from previous codebooks. After initialization, we alterQuantizingnate between updating codes $b_{i,j,m}$ and codebooks $C_{m}$ until the loss function (3) stops improving up to the specified tolerance. Since codes are discrete and codebooks are continuous, and we are optimizing over multiple interacting layers, our approach has three phases, described in Algorithm 1 and detailed below.
3.2 Phase 1: Beam search for codes
First, AQLM updates the codes $b_{i,j,m}$ to minimize the MSE objective (3). Similarly to Babenko & Lempitsky (2014); Martinez et al. (2016, 2018), we reformulate the objective in terms of a fully-connected discrete Markov Random Field (MRF) to take advantage of MRF solvers.
To simplify the derivation, let us first consider a special case of a single output unit ($d_{out}{=}1)$ and a single quantization group (i.e. $g{=}d_{in}$), to get rid of the concatenation operator: $||\mathbf{W}\mathbf{X}-\sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}||_{2}^{2}$. We rewrite this objective by expanding the squared difference:
$||\mathbf{W}\mathbf{X}-\sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}||_{2}^{2}=||\mathbf{W}\mathbf{X}||_{2}^{2}-\\ -2\left\langle\mathbf{W}\mathbf{X}\ \ ,\ \ \sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}\right\rangle_{F}+||\sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}||_{2}^{2}$ (4)
Above, $\langle\cdot,\cdot\rangle_{F}$ denotes a Frobenius inner product of two matrices. Next, let us consider the three components of Eqn. (4) in isolation. First, note that $||\mathbf{W}\mathbf{X}||_{2}^{2}$ is constant in $b$ and can be ignored. The third component can be expanded further into pairwise dot products:
$||\sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}||_{2}^{2}=\sum^{M}_{i=1}\sum^{M}_{j=1}\left\langle C_{i}b_{i}\mathbf{X},C_{j}b_{j}\mathbf{X}\right\rangle_{F}.$ (5)
Note that both the second and third components rely on Frobenius products of $C_{m}b_{m}\mathbf{X}$-like matrices. These matrices can be inconvenient in practice: since $\mathbf{X}\in\mathbb{R}^{d_{in}\times n}$, the size of each matrix will scale with the size of calibration dataset $n$. To circumvent this, we rewrite the products as:
$\left\langle C_{i}b_{i}\mathbf{X},C_{j}b_{j}\mathbf{X}\right\rangle_{F}=\left\langle C_{i}b_{i}\mathbf{X}\mathbf{X}^{T},C_{j}b_{j}\right\rangle_{F}.$ (6)
Thus one can pre-compute $\mathbf{X}\mathbf{X}^{T}\in\mathbb{R}^{d_{in}\times d_{in}}$. We will denote this type of product as $\left\langle\mathbf{A},\mathbf{B}\right\rangle_{\mathbf{X}\mathbf{X}^{T}}\stackrel{{\scriptstyle\mathsf{def}}}{{=}}\left\langle\mathbf{A}\mathbf{X}\mathbf{X}^{T},\mathbf{B}\right\rangle_{F}$ in future derivations. Then, Eqn. (4) becomes:
$||\mathbf{W}\mathbf{X}-\sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}||_{2}^{2}=||\mathbf{W}\mathbf{X}||_{2}^{2}-\\ -2\sum^{M}_{m=1}\left\langle\mathbf{W},C_{m}b_{m}\right\rangle_{\mathbf{XX}^{T}}+\sum^{M}_{i=1}\sum^{M}_{j=1}\left\langle C_{i}b_{i},C_{j}b_{j}\right\rangle_{\mathbf{XX}^{T}}.$ (7)
Finally, we generalize this equation to multiple output units ($d_{out}>1$) and quantization groups ($g{\neq}d_{in}$). For $d_{out}>1$, note that the original objective (3) is additive with respect to output units: thus, we can apply (7) independently to each output dimension and sum up results. To support multiple input groups ($g{\neq}d_{in}$), we can treat each group as a separate codebook where only the codes for the active group are nonzero. Thus, we need to repeat each codebook $d_{in}/g$ times and pad it with zeros according to the active group.
It is now evident that minimizing (4) is equivalent to MAP inference in a Markov Random Field with $\left\langle\mathbf{W},C_{m}b_{m}\right\rangle_{\mathbf{XX}^{T}}$ as unary potentials and $\left\langle C_{i}b_{i},C_{j}b_{j}\right\rangle_{\mathbf{XX}^{T}}$ as pairwise potentials. While finding the exact optimum is infeasible, prior work has shown that this type of MRF can be solved approximately via beam search or ICM (Besag, 1986).
To solve this problem, we chose to adapt a beam search algorithm from Babenko & Lempitsky (2014). This algorithm maintains a beam of $k$ (beam size) best configurations for the codes, starting from the previous solution. On each step, the algorithm attempts to replace one code by trying all $2^{B}k$ alternatives and selecting the $k$ best based on MSE (7).
Since the loss function is additive, changing one code only affects a small subset of loss components. Thus, we can compute the loss function efficiently by starting with a previous loss function (before code replacement), then adding and subtracting the components that changed during this iteration. These few loss components can be computed efficiently by multiplying with $\mathbf{X}\mathbf{X}^{T}$ ahead of beam search. The beam search runs over all $d_{out}$ output units in parallel. This is possible because encoding one output unit does not affect the objective (7) of other units. Note that beam search is not necessarily the best solution to this problem. AQ variants for retrieval (Martinez et al., 2016, 2018) use randomized ICM to find solutions faster. In this study, we chose beam search because it was easier to implement in ML frameworks like PyTorch/JAX.
3.3 Phase 2: Codebook update
In the second phase, we find the optimal codebook vectors $C_{1},...,C_{M}$ that minimize the same squared error as the beam search. If we treat the codes $b$ as constants, minimizing (3) becomes a least squares problem for $C_{m}$. The original AQ algorithm solves this problem in closed form, relying on the fact that each vector dimension can be optimized independently. Our problem is complicated due to the presence of $\mathbf{X}\mathbf{X}^{T}$: the optimal value of one codebook coordinate depends on the values of all others. In principle, we could optimize $C_{m}$ in closed form, but it would require inverting a large matrix, or using iterative least squares solvers (e.g. conjugate gradients) specialized to this problem.
For simplicity, our current implementation defaults to using Adam (Kingma & Ba, 2015) for approximately solving this minimization problem. In practice, this codebook tuning phase takes up a small fraction of the total compute time. We compute the objective as follows:
$||\mathbf{WX}-\widehat{\mathbf{W}}\mathbf{X}||_{2}^{2}=||(\mathbf{W}-\widehat{\mathbf{W}})\mathbf{X}||_{2}^{2}=\\ =\left\langle(\mathbf{W}-\widehat{\mathbf{W}})\mathbf{X}\mathbf{X}^{T},(\mathbf{W}-\widehat{\mathbf{W}})\right\rangle_{F},$ (8)
where $\widehat{\mathbf{W}}$ is the quantized weight matrix from 2, and the $\mathbf{X}\mathbf{X}^{T}$ matrix is pre-computed. We optimize this objective by iterating (non-stochastic) full-batch gradient descent.
For each update phase, our implementation runs 100 Adam steps with learning rate $10^{-4}$. However, we found that the final result is not sensitive to either of these parameters: training with smaller number of steps or learning rate achieves the same loss, but takes longer to converge. In future work, these hyperparameters could be eliminated by switching to dedicated least squares solver for codebooks. Similarly to other algorithms, we also learn per-unit scales $s\in\mathbb{R}^{d_{out}}$ that are initialized as $s_{i}:=||\mathbf{W}_{i}||_{2}$ and updated alongside codebooks via the same optimizer (line 10 in Algorithm 1).
3.4 Phase 3: Fine-tuning for intra-layer cohesion
So far, our algorithm compresses each weight matrix independently of the rest of the model. However, in practice, quantization errors interact differently between matrices. This issue is especially relevant in the case of extreme (2-bit) compression, where quantization errors are larger.
Figure 3: AQLM compressed weight format. Horizontal and vertical axes are input features and output units, respectively. Depth represents the codebook index. Reconstruction procedure, from left to right: i) compressed weight codes ii) zoom-in one weight group, each code is an index in its respective codebook iii) select codes from each codebook iv) add up codes as in (2) v) multiply by scales (one scale per output dimension).
Prior work addresses this issue via quantization-aware training (QAT), e.g. (Gholami et al., 2021). Instead of compressing the entire model in a single pass, they quantize model parameters gradually and train the remaining parameters to compensate for the quantization error. Unfortunately, running QAT in our setting is infeasible, since most modern LLMs are extremely expensive to train or even fine-tune. Thus, most PTQ algorithms for LLMs only adjust model parameters within the same linear layer (Frantar et al., 2022a; Lin et al., 2023; Dettmers et al., 2023b).
Algorithm 1 AQLM: Additive Quantization for LLMs
0:  model, data
[…]
3:    block := model.get_block(i)
4:    $\mathbf{Y}_{block}:=\texttt{block}(\mathbf{X}_{block})$
[…]
7:       $\mathbf{X}:=\texttt{layer\_inputs}(\texttt{layer},\mathbf{X}_{block})$
8:       $C,b,s:=\texttt{initialize}(\mathbf{W})$ // k-means
9:       while  loss improves by at least $\tau$ do
10:          $C,s:=\texttt{train\_Cs\_adam}(\mathbf{XX}^{T},\mathbf{W},C,b,s)$
11:          $b:=\texttt{beam\_search}(\mathbf{XX}^{T},\mathbf{W},C,b,s)$
12:       end while
13:       /* save for fine-tuning */
[…]
15:    end for
[…]
17:    while  loss improves by at least $\tau$ do
18:       $L:=||\texttt{block}(\mathbf{X}_{block})-\mathbf{Y}_{block}||_{2}^{2}$
19:       $\theta:=\texttt{adam}(\theta,\frac{\partial{L}}{\partial{\theta}})$
20:    end while
21:    $\mathbf{Y}_{block}:=\texttt{block}(\mathbf{X}_{block})$
22:  end for
Here, we opt for a middle ground by performing optimization at the level of individual transformer blocks, i.e. groups of 4-8 linear layers [^3] This number depends on factors including the use of gated GLU activations, group query attention and QKV weight merging. that constitute a single multi-head self-attention, followed by a single MLP layer. Having quantized all linear layers within a single transformer block, we fine-tune its remaining parameters to better approximate the original outputs of that transformer block by backpropagating through the weight representation (2).
Concretely, we use the PyTorch autograd engine to differentiate the $||\text{block}(\mathbf{X}_{block})-\mathbf{Y}_{block}||^{2}$ , where $\mathbf{X}_{block}$ are the inputs activations for that transformer block and $\mathbf{Y}_{block}$ are output activations of $\text{block}(\mathbf{X}_{block})$ recorded prior to quantization. We train the codebooks $C_{m}$, scale vectors $s$ and all non-quantized parameters (RMSNorm scales and biases), while keeping the codes $b_{i,j,m}$ frozen. Similarly to Section 3.3, we train these parameters using Adam to minimize the MSE against the original block outputs (prior to quantization). This phase uses the same calibration data as for the individual layer quantization. The full procedure is summarized in Alg. 1.
While fine-tuning blocks is more expensive than individual linear layers, it is still possible to quantize billion-parameter models on a single GPU in reasonable time. Also, since the algorithm only modifies a few trainable parameters, it uses little VRAM for optimizer states. This fine-tuning converges after a few iterations, as it starts from a good initial guess. In practice, fine-tuning transformer layers takes a minority (10-30% or less) of the total calibration time.
Table 1: Evaluation of quantized Llama 2 models for 2-2.8 bits per parameter, with an extra section for higher bitwidth. We report perplexity on WikiText-2 (Merity et al., 2016) & C4 (Raffel et al., 2020) and accuracy for zero-shot tasks. The Average accuracy is the mean of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 2.02 6.59 8.54 65.67 74.76 49.55 63.68 32.76 57.28
QuIP# 2.02 8.22 11.01 62.43 71.38 42.94 55.56 28.84 52.23
AQLM 2.29 6.29 8.11 65.67 74.92 50.88 66.50 34.90 58.57
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 1.97 5.60 7.49 68.82 75.90 53.80 69.28 38.82 61.32
QuIP 2.00 13.48 16.16 52.80 62.02 35.80 45.24 23.46 43.86
QuIP# 2.01 6.06 8.07 63.38 74.76 51.58 64.06 33.96 57.55
AQLM 2.19 5.37 7.16 67.64 77.37 55.03 70.29 38.65 61.80
AQLM 2.53 5.13 6.82 69.77 76.99 56.15 70.33 39.16 62.48
AQLM 2.76 4.94 6.54 68.98 77.58 57.71 72.90 43.60 64.15
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 2.07 3.94 5.72 75.93 80.43 61.79 77.68 47.93 68.75
QuIP 2.01 5.90 8.17 67.48 74.76 50.45 62.16 33.96 57.76
QuIP# 2.01 4.16 6.01 74.11 79.76 60.01 76.85 47.61 67.67
Table 2: Evaluation of quantized Llama 2 models for 3-3.1 bits per parameter, with the same metrics as in Table 1.
[…]
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 3.04 5.46 7.08 66.93 76.88 54.12 68.06 38.40 60.88
GPTQ 3.00 8.06 10.61 59.19 71.49 45.21 58.46 31.06 53.08
SpQR 2.98 6.20 8.20 63.54 74.81 51.85 67.42 37.71 59.07
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.03 4.82 6.37 68.43 77.26 58.30 70.88 42.58 64.49
GPTQ 3.00 5.85 7.86 63.93 76.50 53.47 65.66 38.48 59.61
SpQR 2.98 5.28 7.06 67.48 77.20 56.34 69.78 39.16 61.99
QuIP 3.00 5.12 6.79 69.93 76.88 57.07 70.41 41.47 63.15
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 3.01 3.36 5.17 77.19 81.28 63.23 77.61 50.00 69.86
GPTQ 3.00 4.40 6.26 71.82 78.40 60.00 72.73 44.11 65.41
SpQR 2.98 3.85 5.63 74.66 80.52 61.95 75.93 48.04 68.22
QuIP 3.01 3.87 5.67 74.59 79.98 60.73 73.19 46.33 66.96
4 Experiments
We evaluate the AQLM algorithm in typical scenarios for post-training quantization of modern LLMs. Our evaluation is focused on the Llama 2 model family since it is a popular backbone for fine-tuned models or general LLM applications, e.g. (Dettmers et al., 2023a), and we also present results on Mistral-family models (Jiang et al., 2024). In Section 4.1, we evaluate the full AQ procedure for various Llama 2 models and quantization bit-widths; Section 4.3 presents an ablation analysis for individual AQ components and implementation details.
4.1 Compression quality for modern LLMs
We report perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets. We also measure zero-shot accuracy on WinoGrande (Sakaguchi et al., 2021), PiQA (Tata & Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018) via the LM Eval Harness (Gao et al., 2021). We follow the evaluation setup of GPTQ (Frantar et al., 2022a) and provide configurations for AQLM and baselines in Appendix C.
We consider three main targets in terms of compression ranges: 2-2.8 bits, 3-3.1 bits, and 4-4.1 bits per model parameter. In the results below average bits per parameter takes into account only quantized weights, we do not include parameters kept in floating precision similarly to the related work. The details on the model size estimate are provided in Appendix H. We compare AQ against GPTQ for 3&4 bits (Frantar et al., 2022a), SpQR for 3&4 bits (Dettmers et al., 2023b), QuIP in 2,3 & 4 bits (Chee et al., 2023) and QuIP# for 2&4 bits (Tseng et al., 2024). While GPTQ and SpQR technically support 2-bit quantization, they perform poorly in the 2-3 bit range. For QuIP, our adapted [^4] The official QuIP (non-#) code does not support Llama 2. implementation shows acceptable performance for Llama 2 13B & 70B but performs poorly for the 7B model. We calibrate each algorithm using the subset of RedPajama dataset (Computer, 2023), with a sequence length of 4096.
Table 3: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 2 bits. The table reports perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
8x7B 16 3.46 5.02 75.45 82.37 64.65 83.38 55.80 72.33
AQLM 1.98 4.61 5.75 73.64 79.27 57.91 78.96 48.63 67.68
QuIP# 2.01 4.75 5.89 71.11 79.05 58.23 77.57 45.73 66.34
The exact bit-widths for each method are dictated by parameters such as the number of codebooks and code width. We report results for the $2{-}2.8$ and $3{-}3.1$ bitwidth ranges in Tables 1 and 2, respectively. Additional results for $4-4.1$ bits are deferred to Appendix F.2.
The results show that AQLM outperforms the previous best PTQ algorithms across all settings, often by wide margins, especially at high compression. This holds both in terms of PPL across standard validation sets (Wiki-Text2 and C4), and accuracy across zero-shot tasks. Specifically, we observe the highest accuracy gains in the “extreme” 2-2.1 bits per parameter range, where the deviation from the uncompressed model becomes large for all methods.
Mixtral quantization.
Table 3 presents results on the Mixtral MoE, comparing against QuIP# at 2-bits. (See Appendix F.1 for full results.) AQLM outperforms QuIP# in this case as well. Although the margins are lower compared to Llama 2 models, they are still significant for “harder” tasks, such as Arc Challenge (+3 points).
Pareto optimality of AQLM. The significant error improvements raise the question of choosing the “optimal” model variant to maximize accuracy within a certain memory budget. For this, we follow Dettmers & Zettlemoyer (2022): a quantized model is said to be Pareto-optimal if it maximizes accuracy at the same or lower total size (bytes). Despite rapid progress, prior art methods are not Pareto-optimal at 2-bits: for instance, the previous best 2-bit Llama 2 13B (QuIP#, Table 1) achieves Wiki2 PPL of 6.06, but one can get much lower 5.21 PPL by using a 7B model with 4-bit quantization, which is smaller (see Appendix Table 10).
AQLM compression to strictly 2 bits for the same model is also below Pareto-optimality, as it is outperformed by 4-bit AQLM compression for Llama 2 7B (5.21 vs 5.59). To find the Pareto-optimal quantization bitwidth, we run experiments between 2-3 bits per parameter and report them in Table 1, below horizontal bars. Thus, the Pareto-optimal bitwidth for AQLM appears to be around 2.5 bits per parameter (Table 1), at which point we are comparable to 5-bit AQLM for Llama 2 7B (Appendix Table 10). In turn, the 2.76-bit AQLM on 13B outperforms the uncompressed 7B model. As such, AQLM is the first algorithm to achieve Pareto-optimality at less than 3 bits per parameter.
Table 4: Evaluation of quantized Llama 2 with end-to-end fine-tuning, with the same metrics as in Table 1.
[…]
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 2.02 6.14 8.09 65.67 76.01 51.83 63.43 34.39 58.27
QuIP# 2.02 6.19 8.16 64.96 75.41 51.91 64.96 35.15 58.48
AQLM 2.29 5.92 7.86 63.77 76.93 52.82 66.16 36.95 59.33
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 1.97 5.33 7.19 68.67 76.82 56.31 69.99 40.36 62.43
QuIP# 2.01 5.35 7.20 67.64 77.26 56.04 69.02 39.85 61.96
AQLM 2.19 5.22 6.98 68.27 77.53 57.09 69.78 40.70 62.67
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 2.07 3.83 5.62 74.35 80.90 62.17 74.58 48.98 68.20
QuIP# 2.01 3.91 5.71 74.66 79.54 62.52 77.06 47.61 68.28
4.2 End-to-end fine-tuning experiments
[…]
4.3 Ablation analysis
In Appendix E, we examine key design choices regarding initialization, alternating optimization, the impact of the fine-tuning, and sensitivity to hyperparameters. In brief, we first find that the residual K-means initialization is critical for fast algorithm convergence: when compared with random initialization, it needs significantly fewer training iterations. We also compare different hyperparameter configurations for the same bitwidth, varying the number of codebooks and group size. Second, to validate our calibration fine-tuning procedure, we compare it against 1) no fine-tuning, 2) fine-tuning only of non-linear layers (e.g. RMSNorm) but not of codebook parameters, and 3) fine-tuning only the codebooks (but not other layers). The results, presented in full in Appendix E, show that fine-tuning the codebook parameters has the highest impact on accuracy, by far, while fine-tuning the RMSNorm only has minor impact. This validates our choice of leveraging the calibration set for learned codebooks.
Further, we observe that, increasing the number of sample sequences in the range 128 to 4096 leads to a gradual PPL improvement, but with diminishing returns. This is true for both initial AQLM calibraton and fine-tuning. In this respect, AQLM benefits more from larger calibration sets (similarly to QuIP#), as opposed to direct methods like GPTQ which saturate accuracy at around 256 input sequences. Finally, we investigate various options for investing a given bit budget, comparing e.g. longer codes (e.g. 1x15) vs multiple codebooks with shorter codes (e.g. 2x8).
4.4 Inference Speed
Although our primary objective is to maximize accuracy for a given model size, AQLM can also be practical in terms of inference latency. To demonstrate this, we implemented efficient GPU and CPU kernels for a few hardware-friendly configurations of AQLM. The results can be found in Table 5. For GPU inference, we targeted quantized Llama 2 models with 16-bit codebooks, corresponding to 2.07 bits for Llama 2 70B, 2.19 bits for 13B, and 2.29 bits for 7B models (see Table 1, 4), as well as a 2x8-bit codebook model with perplexity 6.57 on Wiki2(see Table 12). For each model we benchmark the matrix-vector multiplication subroutine performance on a standard layer. The results show that AQLM can execute at speeds comparable to or better than FP16. End-to-end generative numbers with HuggingFace integration can be found in Appendix I: for instance, we can achieve ${\approx}14$ tokens/s on Llama 2 70B in this setting. We observe that multiple smaller codebooks allow efficient GPU cache utilization, leading to greater speedup, at the price of slightly lower accuracy.
Table 5: Speed of the FP16 gate_proj layer matrix-vector multiplication in PyTorch, and relative AQLM speedups.
Llama 2 7B 13B 70B
2 bit speedup over FP16 on Nvidia RTX 3090 GPU
Original (float16) 129 $\mu$s 190 $\mu$s 578 $\mu$s
AQLM (Table 1) x1.31 x1.20 x1.20
AQLM (2×2$\times$2 ×8-bit) x1.57 x1.82 x3.05
2 bit speedup over FP32 on Intel i9 CPU, 8 cores
Original (float32) 1.83 ms 3.12 ms 11.31 ms
AQLM (2×2$\times$2 ×8-bit) x2.75 x3.54 x3.69
AQLM (4×4$\times$4 ×8-bit) x2.55 x3.02 x4.07
AQLM (8×8$\times$8 ×8-bit) x2.29 x2.68 x4.03
Next, we explore how to leverage AQLM to accelerate CPU inference. As discussed in Section 2.2, additive quantization can compute dot products efficiently if the codebook size is small. One way to achieve it for AQLM is to replace each 16-bit codebook with a number of smaller 8-bit ones. This leads to higher quantization error, but still outperforms the baselines in terms of accuracy (see Appendix Table 9). The results in Table 5 show that this also allows for up to 4x faster inference relative to FP32 on CPU.
5 Conclusion and Future Work
We presented AQLM, a new form of additive quantization (AQ) targeting LLM compression, which significantly improved the state-of-the-art results for LLM quantization in the regime of 2 and 3 bits per weight. In terms of limitations, AQLM is more computationally-expensive than direct post-training quantization methods, such as RTN or GPTQ, specifically because of the use of a more complex coding representation. Yet, despite the more sophisticated encoding and decoding, we have shown AQLM lends itself to efficient implementation on both CPU and GPU. Overall, we find it remarkable that, using AQLM, massive LLMs can be executed accurately and efficiently using little memory.
While AQLM already achieves substantial improvements in low-bit quantization, there are several promising directions for further improvement that we did not explore in this work. One such direction is better fine-tuning strategies. In Section 4.2 we found that better fine-tuning algorithms (Tseng et al., 2024; Malinovskii et al., 2024) can significantly improve quantized model accuracy. We believe that AQLM can benefit from a more systematic exploration of fine-tuning algorithms in future work. Another promising direction is generalizing AQLM to other quantization scenarios. While our work is focused around LLM quantization, the underlying algorithm can potentially be adapted to other problems, e.g. quantizing computer vision models, compressing LLM attention caches for long sequences, and others.
Acknowledgements
Authors would like to thank Ruslan Svirschevski for his help in solving technical issues with AQLM and baselines. We also thank Tim Dettmers for helpful discussions on the structure of weights in modern LLMs and size-accuracy trade-offs. The authors would also like to thank Daniil Pavlov for his assistance with CPU benchmarking. Finally, authors would like to thank the communities of ML enthusiasts known as LocalLLaMA [^5] https://www.reddit.com/r/LocalLLaMA/ and Petals community on discord [^6] https://github.com/bigscience-workshop/petals/ for the crowd wisdom about running LLMs on consumer devices. Egiazarian Vage and Denis Kuznedelev and Andrei Panferov were supported by the grant for research centers in the field of AI provided by the Analytical Center for the Government of the Russian Federation (ACRF) in accordance with the agreement on the provision of subsidies (identifier of the agreement  [^00] D730321P5Q0002) and the agreement with HSE University No. 70-2021-00139
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A End-to-end fine-tuning
The block-wise finetuning procedure, introduced in 3.4, considerably improves performance of compressed models. However, block-wise finetuning optimizes the loss only at the level of a current transformer block and is agnostic of the actual task of interest. To minimize the target loss, one can run backpropagation through the whole model and directly optimize all trainable parameters to minimize a model-level objective function.
This allows to search for globally optimal parameters, as opposed to sequentially selected ones, during block-wise finetuning.
One can minimize the error between the quantized model and the floating-point model on some calibration set. The parameters being optimized (namely the codebooks, scales and the non-quantized parameters) typically constitute a small fraction of the total number of parameters in the original model. Therefore, the proposed distillation method resembles parameter-efficient finetuning (PEFT) in both optimization and memory footprint.
To transfer the knowledge from the original model to the quantized one, we adopt Knowledge Distillation (Hinton et al., 2015) where the student model is taught to mimic the output of a teacher given the same input. We follow the setup from QuIP# (Tseng et al., 2024) that uses KL divergence between the outputs of teacher and student models:
$\mathcal{L}=\frac{1}{N}\sum_{i=0}^{N-1}D_{KL}(p_{s}(\mathbf{x}_{i}),p_{t}(\mathbf{x}_{i}))$ (9)
Above $D_{KL}$ is the Kullback–Leibler divergence and $p_{s}$, $p_{t}$ are the student and teacher probabilities given input sequence $\mathbf{x}_{i}$.
Despite its simplicity, this fine-tuning procedure often significantly improves performance of the compressed model.
We fine-tune all models on $4{-}16$M training tokens: $1{-}4$k sequences of length 4k for Llama 2 models (Touvron et al., 2023) and 512 sequences of length 8k for Mixtral (Jiang et al., 2024). We fine-tune on the same data as during initial calibration (i.e. samples from RedPajama (Computer, 2023)) and use Adam (Kingma & Ba, 2015) optimizer with constant learning rate $10^{-5}$ without weight decay. Batch size is set to $8{-}16$ sequences. A single epoch of fine-tuning turns out to be sufficient, and longer training leads to marginal improvements.
Appendix B Code reproducibility
We share the code for our method in the GitHub repository https://github.com/Vahe1994/AQLM/tree/AQLM_camera_ready. The hyperparameters for our experimental setup are discussed in Appendix C.
Appendix C Experimental Configurations
Hardware. In all of our experiments, we used either Nvidia A100 or H100. The number of GPUs varied from 1 to 8. We used activation offloading to lower pick memory usage. To evaluate inference speed on GPU we used consumer-grade GPU Nvidia 3090 and for CPU setup we used Intel core i9 13900k.
Calibration set. All methods were calibrated on a slice of RedPajama-v1 dataset (Computer, 2023) for both Llama and Mistral/Mixtral family models. We used the same context length as models were trained on, for Llama 2 4096 and for Mistral/Mixtral 8192.
For Llama 2 experiments, we used 8M tokens as a calibration set for SpQR, GPTQ, and AQLM. Quip, however, was calibrated on 4M tokens due to OOM errors when trying to use more samples. Taking into account the fact that after 2M tokens improvement of methods results is fairly small we chose to report these numbers as is. For Quip#, we used Llama 2 and Mistral’s quantized models provided by authors in their GitHub repository. To the best of our knowledge, they used 6k samples for calibration with a context length of 4096/8192.
For Mixtral we calibrated both our method and QUIP# on 8M tokens with context length 8192.
Hyperparameters.
For GPTQ for both 3 and 4 bits we used a standard set of parameters without grouping and with permutation order act_order.
SpQR method was evaluated with base 2 and 3 bit-width with group size of 16 and 3 bits for zeros and scales. Outliers rate was chosen such that average bit will be close to 3 and 4 bits respectively.
Quip was adapted to work on the Llama family and was calibrated with 1024 samples and 4096 context length.
Quip# For Llama 2 and Mistral models we used the officially published quantized models. For Mixtral we adapted the code to work with the model’s architecture and quantized it with the recommended set of parameters. For both AQLM and QuIP#, we don’t quantize gate linear layer in Mixtral, because it contains relatively small amount of paramters and have severe impact on performance.
AQLM For to get 2, 3, 4 bits: we used 1 codebook size of $2^{15}$ or $2^{16}$, with groups of 8 for 2 bits. For 3 bits we used 2 codebooks size of $2^{12}$ with groups of 8. Finally for 4 bits we used 2 codebooks size of $2^{15}$ or $2^{16}$ with groups of 8.
Both for finetuning 3.4 and codebooks update 3.3 we used Adam optimizer (Kingma & Ba, 2015) with learning rate of $10^{-4}$, $\beta_{1}=0.90$ and $\beta_{2}=0.95$. We used early stopping both for the finetuning phase and for the codebook optimization phase, by stopping when the least square error not decreasing more than some threshold. In our experiments the threshold varies between $10^{-2}$ and $10^{-3}$.
Hyperparameters for end-end fine-tuning discussed at the end of Appendix A.
Appendix D Quantization time
AQLM quantization takes considerably longer to calibrate than simpler quantization methods such as RTN or GPTQ. This only impacts quantization time, not inference time.
Quantizing a 7B model with default configuration takes about 1 day on a single A100 gpu. Similarly, quantizing a 70B model on a single GPU would take 10-14 days. However, the procedure can be parallelized across multiple GPU: 7B quantization takes  14h on 2 GPUs, and 70B quantization takes 3-4 days on 8 GPUs.
Full model fine-tuning with default configuration for 7B model would take 3-6 hours on four A100 , for 13B 10-16 hours on four A100, and for 70B 1-2 days on 8 A100.
Finally, the quantization time is dependent on the quantization configuration and its hyperparameters. Tweaking these parameters, e.g. by reducing the number of beams, can achieve notable speedups of 2-4x during quantization, but at the cost of lower model accuracy.
Table 6: Evaluation of quantized Llama 2 end-to-end fine-tuned models for 3-3.1 bits per parameter, with the same metrics as in Table 1.
[…]
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 3.04 5.38 7.01 65.35 77.31 55.49 66.79 38.48 60.68
QuIP# 3.04 5.41 7.04 66.85 77.31 55.32 68.43 38.99 61.38
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.03 4.78 6.33 68.75 78.45 58.54 72.94 42.75 64.29
QuIP# 3.01 4.78 6.35 68.03 77.86 57.56 72.18 41.38 63.40
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 3.01 3.36 5.17 75.30 80.69 63.48 77.99 50.26 69.54
QuIP# 3.00 3.35 5.15 76.40 81.45 63.53 77.53 50.77 69.94
Appendix E Ablation analysis
The AQLM algorithm makes several design choices that need to be validated separately: initialization, alternating optimization, the fine-tuning protocol, and the choice of hyperparameters. Here, we study how each of these components affect results.
Initialization. As discussed in Section 3, we initialize AQLM with residual K-means to obtain a good initial guess for both codes and codebooks. That is, we run K-means for the weight matrix, then subtract the nearest cluster from each weight, and run K-means again M times. A simple baseline would be to initialize all codes uniformly at random. We compare the two initialization strategies for the problem of quantizing a single linear layer within Llama 2 70B model to 3 bits per parameter. We quantize groups of 8 consecutive weights using 2 codebooks, 12 bit each. Each codebook contains ${2^{12}}$ learnable values. As we can see in Figure 4, AQLM with K-means initialization needs significantly fewer training iterations to achieve the desired loss. The difference is so drastic that we expect that running AQLM with a random initialization would require extremely high runtimes to accurately quantize the largest models.
Figure 4: MSE loss learning curves of AQLM trained on the self attention q_proj linear layer of 10-th block in the Llama 2 70B model.
Fine-tuning. Next, we validate the fine-tuning procedure. We compare the full block fine-tuning (default) against three alternatives: i) no fine-tuning at all, ii) fine-tuning only non-linear layers (i.e. RMSNorm), but not the AQ parameters, and iii) fine-tuning only the AQ parameters, but not the non-linear layers. Table 7 summarizes our results: fine-tuning the entire model or only AQ parameters achieves competitive performance, while training only RMSNorm scales is comparable to not fine-tuning at all. We attribute these observations to the fact that over 99% of quantized layer parameters are contained in AQ codebooks $C_{m}$, whereas the remaining parameters are small 1-dimensional tensors. This validates the use of the AQ approach, as many competing algorithms do not have learnable per-layer codebooks. Notably, QuIP# uses a shared fixed lattice instead. We also note that, even without fine-tuning, AQLM is competitive to previous state-of-the-art results.
Table 7: Ablation analysis of AQLM with different fine-tuning restrictions on Llama-2 7B model at 2.02 bit width.
Name Wiki2$\downarrow$ C4$\downarrow$
w/o 8.18 10.59
RMSnorm 8.31 10.46
AQ params 6.92 8.85
Full 6.93 8.84
Number of samples. We verify our choice of calibration hyperparameters. Traditionally, most PTQ algorithms use several hundred calibration sequences (e.g. Frantar et al. (2022a) has 128). In our experiments, we evaluate both AQLM and baselines with additional calibration data. Our original motivation for that was to avoid potential overfitting when fine-tuning entire transformer blocks. To test this assumption, we run our algorithm with different calibration set sizes, varying from 128 to 4096 sequences. For each size, we report the average perplexity on WikiText-2 over 3 runs, along with standard deviations. The results in Table 8 demonstrate that increasing the number of samples leads to gradual reduction in perplexity with seemingly diminishing returns. Since the perplexity is still monotonically improving from 128 to 4096 samples, it is possible that larger sample sizes would yield further improvements.
Number of codebooks vs groups. Finally, we conducted an additional set of experiments on LLama 2 7B models to see perplexity dependence on simultaneous change on WikiText-2 of both codebooks and groups keeping compression rate fixed to 2bits. We present both AQLM with and without end-to-end fine-tuning in Table 9.
Table 8: WikiText-2 PPL as a function of calibration set size for Llama 2 (7B) quantized to 2.3 bits with AQLM, averaged over 3 runs. SD stands for adjusted standard deviation.
# of samples Average PPL SD
128 6.994 0.127
256 6.584 0.031
512 6.455 0.005
1024 6.353 0.008
2048 6.297 0.018
4096 6.267 0.005
Table 9: WikiText-2 PPL as a function of from groups and number of codebook for Llama 2 (7B) quantized with approximately 2 bits quantization.
Method Setup Average PPL
AQLM 2x8gs8 7.6107
4x8gs16 8.1394
8x8gs32 7.3755
15x8gs64 7.8459
AQLM 2x8gs8 6.5746
8x8gs32 6.6126
15x8gs64 6.6602
Appendix F Additional experiments
In this section we report additional experimental results for Mixtral(Jiang et al., 2024), Mistral7B(Jiang et al., 2023) and LLama 2 model.
Table 10: Evaluation of quantized Llama 2 models for 4+ bits per parameter. The table reports perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 4.04 5.21 6.75 67.32 78.24 55.99 70.16 41.04 62.55
GPTQ 4.00 5.49 7.20 68.19 76.61 55.44 66.20 36.77 60.64
SpQR 3.98 5.28 6.87 66.93 78.35 56.10 69.11 39.68 62.17
QuIP# 4.02 5.29 6.86 66.85 77.91 55.78 68.06 39.68 61.66
AQLM 5.02 5.16 6.68 67.40 78.29 56.53 68.94 39.93 62.22
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.94 4.65 6.14 69.85 78.35 59.27 73.32 44.80 65.12
GPTQ 4 4.78 6.34 70.01 77.75 58.67 70.45 42.49 63.87
SpQR 3.98 4.69 6.20 69.69 78.45 59.25 71.21 44.52 64.42
QuIP 4.00 4.76 6.29 69.69 79.00 58.91 73.27 44.88 65.15
QuIP# 4.01 4.68 6.20 69.38 77.91 58.86 73.74 44.63 64.90
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 4.14 3.19 5.03 76.48 81.50 63.69 77.31 50.68 69.93
GPTQ 4.00 3.35 5.15 75.61 81.23 63.47 76.81 49.15 69.25
SpQR 3.97 3.25 5.07 76.01 81.28 63.71 77.36 49.15 69.50
QuIP 4.00 3.58 5.38 76.01 80.25 61.97 74.28 47.01 67.90
QuIP# 4.01 3.22 5.05 76.80 81.45 63.51 78.37 50.85 70.20
AQLM 3.82 3.21 5.03 76.32 80.90 63.69 77.61 50.34 69.77
F.1 Mixtral
We report the results for Mixtral(Jiang et al., 2024) MoE-type model for 3 and 4 bits in Table 11. In the 4 bit case, performance of QuIP# and AQLM are very similar across all metrics and close to uncompressed FP16 model.
Table 11: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 3 and 4 bits per parameter. The table reports perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. The primary metrics are Wiki2 (PPL, lower is better), C4 (PPL, lower is better) and Average accuracy (percentage, higher is better).
[…]
3-bit 16.00 3.46 5.02 75.45 82.37 64.65 83.38 55.80 72.33
AQLM 3.02 3.79 5.17 75.45 81.61 63.25 81.90 53.92 71.23
4-bit 16.00 3.46 5.02 75.45 82.37 64.65 83.38 55.80 72.33
AQLM 3.915 3.57 5.07 74.82 81.99 64.23 83.12 54.61 71.75
QuIP# 4.000 3.60 5.08 76.56 81.99 63.92 82.62 54.78 71.97
F.2 Llama 2
We show results for 4 bit quantization of the Llama 2 models in Table 10. We can see that AQLM outperforms other methods in terms of perplexity and has the best or close to the best results. We also report results of perplexity for our quantized 2x8 codebooks models in Table 12.
Table 12: Evaluation of quantized Llama 2 for 2x8groupsize8 codebooks models. We report perplexity on WikiText-2 (Merity et al., 2016) & C4 (Raffel et al., 2020) and accuracy for zero-shot tasks. The Average accuracy is the mean of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
7B 16 5.12 6.63 67.25 78.40 56.67 69.36 39.51 62.24
AQLM 2 7.61 9.68 62.27 71.87 46.41 61.03 30.03 54.32
AQLM 2 6.57 8.60 63.22 74.54 50.08 61.28 31.83 56.19
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 2 6.54 8.77 55.96 71.06 48.29 62.50 31.40 53.84
AQLM 2 5.63 7.55 6385 77.04 54.19 67.85 37.20 60.03
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 2 4.21 5.99 73.48 79.54 61.29 74.49 46.84 67.13
Table 13: Evaluation of quantized Mistral7B (Jiang et al., 2023) models for 2, 3 and 4 bits per parameter: perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
2-bit 16.00 4.77 5.71 73.64 80.47 61.15 78.87 49.23 68.67
AQLM 2.01 6.32 6.93 68.75 76.01 52.13 73.65 40.44 62.17
QuIP# 2.01 6.02 6.84 69.30 76.71 52.95 72.14 39.76 62.20
AQLM 2.01 5.76 6.60 68.67 77.64 56.44 73.32 42.66 63.75
3-bit 16.00 4.77 5.71 73.64 80.47 61.15 78.87 49.23 68.67
AQLM 3.04 5.02 5.93 73.24 79.22 59.31 78.28 46.76 67.36
AQLM 3.04 5.12 6.09 72.85 79.05 59.92 77.57 48.12 67.50
4-bit 16.00 4.77 5.71 73.64 80.47 61.15 78.87 49.23 68.67
AQLM 4.02 4.89 5.81 73.80 79.71 60.27 77.86 48.21 67.97
QuIP# 4.01 4.85 5.79 73.95 80.41 60.62 78.96 49.40 68.67
F.3 Mistral
Finally, we evaluate AQLM and QuIP# quantization on Mistral 7b (Jiang et al., 2023) model for 3 and 4 bits in Table 13. In 2 bits, QuIP# slightly outperform AQLM on most benchmarks. And for 4 bits setup results are very close across the board.
Appendix G Pareto optimality
We visualize WikiText-2 perplexity of Llama-2 7B, 13B, 70B models quantized with AQLM and QuIP# as plotted against quantized weight size in bytes and report it in Figure 5. Our method outperforms QuIP# in terms of perplexity in WikiText-2 across all model sizes.
Additionally, in Figure 6, we show perplexity on WikiText-2 for AQLM method against size of quantized parameters. We can notice that starting around 3.7 GiB of quantized weights, which correspond to 2.5 bits compression on Llama 2 13B model, it is more advantageous to compress 13B model rather 7B model at the same model size in bytes.
Figure 5: Comparison of AQLM relative to QuIP# on Llama 2 7B, 13B, and 70B models.
Figure 6: Model optimality for AQLM on Llama 2 7, 13, and 70B models.
Appendix H Estimating model size
In this section, we describe how to estimate the size of the quantized model for a given codebook configuration. The total cost of storing quantized weight comprises the codebooks, codes and per-unit scales. Specifically for a weight with input dimension $d_{in}$, output dimension $d_{out}$, group size $g$, $M$ codebooks corresponding to $B$-bit codes, the total amount of memory required is (assuming that codebooks and scales are stored in half precision):
  • codebooks: $g\cdot M\cdot 2^{B}\cdot 16$
  • codes: $d_{out}\cdot(d_{in}/g)\cdot B$
  • scales: $d_{out}\cdot 16$
Therefore, the average bits per parameter can be computed as follows:
[…]
For example, for mlp.gate_proj layer of Llama 2 70B model with $d_{in}=8192$, $d_{out}=28672$, quantization with group size $8$, two $8$-bit codebooks the formula above yields $2.002$ bits per parameter. Typically, storage cost is dominated by the codes, whereas codebooks and scales induce small memory overhead.
Appendix I End-to-End Inference Speed
Table 14: Text generation speed benchmark.
Llama 2 7B 13B 70B
Inference on Nvidia RTX 3090 GPU, tok/s
Original (float16) 54.2 29.5 5.8
AQLM (1×1$\times$1 ×16-bit) 65.3 34.1 6.7
AQLM (2×2$\times$2 ×8-bit) 114.1 68.1 14.3
Inference on Intel i9 CPU, 8 cores, tok/s
Original (float32) 3.106 1.596 0.297
AQLM (2×2$\times$2 ×8-bit) 6.961 4.180 0.966
AQLM (4×4$\times$4 ×8-bit) 6.837 4.004 0.948
AQLM (8×8$\times$8 ×8-bit) 5.319 3.193 0.775
For quantized Llama 2 models, setup described in Section 4.4, we measure the time it takes to generate 128 tokens from scratch, performed on compiled computational graphs, with batch size 1, and report the average number of generated tokens per second on a single 24GB RTX 3090 GPU, as well as Intel i9 CPU, in Table 14. Perplexity on WikiText-2 on these configurations presented at the Table 9
Appendix J Codebook and codes distribution
The proposed AQLM quantization method allows for large freedom in the choice of quantization lattice and ability to represent different weight distribution. To understand how do the learned codes and codebooks look like, we visualize the distribution of codes (how frequently given codebook vector is chosen) and the learned codebooks. Below on Figure 7 we provide a cumulative probability plot of leaned codes and two leading principal codebook components for a specific layer. One can observe that codes distribution is close to uniform. Its entropy equals 15.91 bits per code, which is close to the maximum possible entropy of 16 bits (for a 16-bit codebook) for the uniform distribution. Codebook vectors are concentrated in some ball. This pattern is pertinent to all linear projections inside transformer blocks.
Figure 7: Visualization of learned codes and codebooks in layers.5.self_attn.q_proj linear projection.
(Left) Codes distribution. (Right) Two leading principal components of codebook.
Appendix K Evaluation on MMLU and GSM8k
While measurement of perplexity on WikiText-2 and C4 together with zero-shot accuracy on subset of simple 0-shot tasks from LM Eval Harness (Gao et al., 2021) is an established benchmark for evaluation of performance of compressed models, it may be not exhaustive enough for many real-world cases. While the complete and exhaustive evaluation of LLM abilities is still an open question, we evaluate our AQLM models and QuIP# on MMLU (Hendrycks et al., 2020) benchmark that involves problems from 57 different domains, such as humanities, social sciences, physics, e.t.c, and GSM8k (Cobbe et al., 2021) to assess the performance of quantized models on more complex and challenging tasks, requiring reasoning to get the correct answer. Below we consider AQLM and QuIP# after end-to-end finetuning, i.e. the best performing quantized models. We observed that relative decrease on performance on these tasks is higher compared to the standard evaluation. Fine-tuned AQML and QuIP# yield very similar performance on these benchmarks.
Table 15: Evaluation of quantized Llama 2 models for 2-2.1 bits per parameter on MMLU and GSM8k.
corresponds to end-to-end finetuning
Size Method Avg bits MMLU (5-shot) GSM8k (8-shot)
7B 16 45.9 14.6
QuIP# 2.02 36.8 6.2
AQLM 2.02 38.5 5.3
13B 16 55.2 24.3
QuIP# 2.01 50.0 14.0
AQLM 1.97 48.8 13.8
70B 16 68.8 56.3
QuIP# 2.01 65.3 46.4
AQLM 2.07 65.3 47.9
Appendix L Block-wise tuning for scalar quantization
The block-wise procedure introduced in our work is quite general and can be applied to scalar quantization as well. Specifically, operations with quantized weights are differentiable with respect to quantization scales kept in original precision. Therefore, scales can be tuned in the same way as AQLM codebooks. We observed that tuning significantly improves the quality of GPTQ at low bit widths. However, the resulting quality is still far below AQLM at similar bit-widths.
Table 16: Evaluation of AQLM and GPTQ quantization after block tuning for Llama 2 models with 2-2.1 bits per parameter.
Size Method Avg bits Wiki2$\downarrow$ C4$\downarrow$
7B 16 5.12 6.63
GPTQ 2.14 16.77 17.53
AQLM 2.02 6.64 8.56